# An algebraic characterization for spectral radii of trees

### v2

@@ Abstract
There have been various results on spectral radii of trees, studying the relations between spectral radius $\lambda(T)$ of a tree $T$ and some other parameters of $T$. In this article, we provide an algebraic characterization for the spectral radii of trees, and apply this new result to study the existence of trees with given spectral radius and maximum degree.
For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set defined recursively: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,1\leq i\leq s\}$, if $\beta:=\alpha-\sum\limits_{i=1}^{s}q_{i}^{-1}\geq 0$ and $s\leq r-\left\lceil{\frac{\beta}{\beta+1}}\right\rceil$, then $\beta\in{\mathscr{W}}_{r}(\alpha)$. We first show that for any positive integer $r$ and real number $\alpha\geq 1$, there exists a tree $T$ with maximum degree $\Delta(T)\leq r$ and spectral radius $\lambda(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$. It follows that the set of spectral radii of non-trivial trees is exactly the set of positive numbers $\alpha$ such that $0\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$. Applying this conclusion, we then prove that for any positive integers $r\geq 2$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
@@ 1 Introduction
All graphs considered in this article are simple and undirected. For a graph $G$, let $V(G)$, $E(G)$ and $\Delta(G)$ denote its vertex set, edge set and maximum degree, respectively. For any $u\in V(G)$, let $N_{G}(u)$ (or simply $N(u)$) denote the set of neighbours of $u$ in $G$, and let $d_{G}(u)$ or (simply $d(u)$) be the size of $N_{G}(u)$, called the degree of $u$ in $G$. Let $L(G)$ denote the set of vertices $v\in V(G)$ with $d_{G}(v)=1$.
Let ${\mathbb{N}}^{+}$ denote the set of positive integers and for any $k\in{\mathbb{N}}^{+}$, let $\llbracket k\rrbracket=\{i\in{\mathbb{N}}^{+}:i\leq k\}$. For a graph $G=(V,E)$, where $V=\{v_{i}:i\in\llbracket k\rrbracket\}$, its adjacency matrix $A(G)$ according to the vertex ordering $v_{1},v_{2},\ldots,v_{n}$ is defined to be the $0-1$ matrix $(a_{i,j})_{n\times n}$, where $a_{i,j}=1$ if and only if $v_{i}v_{j}\in E$. The characteristic polynomial of $G$, denoted by $P(G,\lambda)$, is defined to be the function $det(\lambda I_{n}-A(G))$, where $I_{n}$ is the identical matrix of size $n$. It is known that $P(G,\lambda)$ is independent of the vertex ordering $v_{1},v_{2},\dots,v_{n}$, and the roots of the equation $P(G,\lambda)=0$, called the eigenvalues of $G$, are real numbers (see [6]), and the largest eigenvalue of $G$, denoted by $\lambda(G)$, is called the spectral radius of $G$.
@@
The spectral radii of graphs, which have been studied extensively (see a survey [26] due to Stevanovíc), plays a crucial role in spectral graph theory, with applications in spanning combinatorics, network theory, and even theoretical physics (see [1, 2, 3, 4, 9, 12, 13, 22, 23, 26]). The distribution of spectral radii among different graphs is a topic of ongoing research with connections to various graph properties and invariants (see [7, 12, 13, 16, 20, 22, 23]).
@@
In this article, we focus on the study of spectral radii of trees. There are various known results on the relation between $\lambda(T)$ and other parameters of a tree $T$. Guo and Tan [11] found an upper bound of $\lambda(T)$ in terms of the order and matching number of $T$. Lovász and Pelikán [15] determined the maximal and minimal value of $\lambda(T)$ among all trees $T$ of order $n$, Simić and Tos̆ić [24] found the tree that has the largest spectral radius among all trees with given order and maximum degree, and Stevanović [25] provided a relation between $\lambda(T)$ and $\Delta(T)$: for any tree $T$ with $|V(T)|\geq 3$,
$\frac{1}{4}\lambda(T)^{2}+1<\Delta(T)\leq\lambda(T)^{2}.$ (1)
In this article, we will first provide a characterization on $r$ and $\alpha$ for the existence of trees $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$. Let $\Lambda$ denote the set of $\lambda(T)$ for all non-trivial trees $T$ (i.e., trees of order at least $2$). Clearly, $\alpha\geq 1$ for each $\alpha\in\Lambda$. Let ${\mathbb{R}}^{+}$ denote the set of positive real numbers, and for any $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of non-negative real numbers defined below:
@@
It can be verified that ${\mathscr{W}}_{1}(1)=\{0,1\}$, ${\mathscr{W}}_{1}(\alpha)=\{\alpha\}$ for all $\alpha\in{\mathbb{R}}^{+}$ with $\alpha\neq 1$, ${\mathscr{W}}_{2}(\sqrt{2})=\left\{0,\sqrt{2},(\sqrt{2})^{-1}\right\}$ and ${\mathscr{W}}_{2}(2)=\big{\{}\frac{i+1}{i}:i\geq 1\big{\}}$. Clearly, the property ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ holds for any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$.
Our first aim in this article is to establish a connection between the existence of a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ and the existence of 00 in the set ${\mathscr{W}}_{r}(\alpha)$.
Theorem 1.1.
For any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$, there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$.
A characterization of the set $\Lambda$ follows from Theorem 1.1 and (1) directly.
Corollary 1.2.
$\Lambda=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)\big{\}}$.
@@
Corollary 1.3.
$\Lambda=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}(\alpha)\big{\}}$.
The conclusion of (1) tells that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$ is a necessary condition for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$. We wonder if this condition is also sufficient.
Conjecture 1.4.
Let $\alpha\in\Lambda\setminus\{1\}$. If $r$ is an integer such that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, then there exists a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
Our second aim is to apply Theorem 1.1 to prove Conjecture 1.4 for the case $\alpha=\sqrt{k}$ for each $k\in{\mathbb{N}}^{+}\setminus\{1\}$. Certainly, it includes the case $\alpha=k$ for each $k\in{\mathbb{N}}^{+}\setminus\{1\}$.
Theorem 1.5.
For any $k,r\in{\mathbb{N}}^{+}\setminus\{1\}$, there exists a tree $T$ with $\lambda(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
The remainder of this article is structured as follows. In Sections 2 and 3, we prove Theorem 1.1’s necessity and sufficiency, respectively. In Section 4, we complete the proof of Theorem 1.5. Finally, in Section 5, we provide some open problems arising from our work.
2 The necessity of Theorem 1.1
The Perron–Frobenius theorem, proved by Oskar Perron [19] in 1907 and Georg Frobenius [10] in 1912, is an important result in the study of eigenvalues and eigenvectors of square matrices. Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectral radius $\lambda(G)$ of a connected graph $G$ is obtained. Let $\lambda_{i}(G)$ denote the $i$-th largest eigenvalue of $G$. Then, $\lambda(G)=\lambda_{1}(G)$. We say $G$ is non-trivial if $|V(G)|\geq 2$.
Theorem 2.1 (Perron–Frobenius theorem).
@@
Applying Theorem 2.1, the next conclusion follows directly.
Proposition 2.2.
Let $G$ be a connected non-trivial graph and $\alpha$ be a positive number. Then, $\lambda(G)=\alpha$ if and only if there is a mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
@@
Note that for any mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$, the condition of (2) holds if and only if $\phi$ is an eigenvector of $A(G)$ with respect to $\lambda(G)$. Now we are going to prove the following lemma by which the necessity of Theorem 1.1 follows directly.
Lemma 2.3.
Let $T$ be a non-trivial tree and $\alpha=\lambda(T)$, and let $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ be an eigenvector of $A(T)$ with respect to $\alpha$. Then both $\frac{\phi(u_{1})}{\phi(u_{2})}$ and $\frac{\phi(u_{2})}{\phi(u_{1})}$ belong to ${\mathscr{W}}_{r}(\alpha)$ for each $u_{1}u_{2}\in E(T)$, where $r=\Delta(T)$.
Proof. Since $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is an eigenvector of $A(T)$ with respect to $\alpha=\lambda(T)$, (2) holds for each $u\in V(T)$. Let’s first prove the following claim.
@@
Proof of the necessity of Theorem 1.1. We follow the conditions in Lemma 2.3. Take any edge $u_{1}u_{2}\in E(T)$, where $u_{1}$ is a leaf of $T$. We have $\alpha\phi(u_{1})=\phi(u_{2})$ by (2), implying that $\alpha^{-1}=\frac{\phi(u_{2})}{\phi(u_{1})}\in{\mathscr{W}}_{r}(\alpha)$ by Lemma 2.3. Thus, $0=\alpha-\frac{1}{\alpha^{-1}}\in{\mathscr{W}}_{r}(\alpha)$ by definition.
Hence the result holds. $\Box$
3 The sufficiency of Theorem 1.1
@@
By Lemma 3.1, if $0\in{\mathscr{W}}_{r}(\alpha)$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$, and there exists $(T,\omega)$ in ${\mathscr{T}}_{r,\alpha}(0)$.
Lemma 3.2.
For any $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\lambda(T)=\alpha$.
Proof. By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$. It suffices to show that $\lambda(T)=\alpha$.
By definition again, $T$ is a directed tree with a unique source $z$ and $\omega(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)$. Let $T_{1},T_{2},\dots,T_{s}$ be the components of $T-z$, where $s=od_{T}(z)$. Clearly, $1\leq s\leq r$. For each $i\in\llbracket s\rrbracket$, $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(\beta_{i})$, where $\omega_{i}$ is the restriction of $\omega$ to $V(T_{i})$, $\beta_{i}=\omega(z_{i})$ and $z_{i}$ is the vertex in $T_{i}$ with $z_{i}\in N^{+}_{T}(z)$. Then $z_{i}$ the unique source of $T_{i}$ for each $i\in\llbracket s\rrbracket$ and
@@
In order to show that $\lambda(T)=\alpha$, by Proposition 2.2, it suffices to prove the following Claim.
@@
where the last step follows from (5). Thus, $\phi$ satisfies condition (2) for each $u\in V(T)$, implying that $\lambda(T)=\alpha$ by Proposition 2.2. $\Box$
By Lemmas 3.1 and 3.2, if $0\in{\mathscr{W}}_{r}(\alpha)$, then there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$, and thus the sufficiency of Theorem 1.1 holds.
Corollary 3.3.
If there exist a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r\rrbracket\big{\}}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
Proof. Assume that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, where $q_{i}\in{\mathscr{W}}_{r}(\alpha)$ and $q_{i}>0$ for each $i\in\llbracket r\rrbracket$. By Lemma 3.1, there exists $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket r\rrbracket$. Assume that $z_{i}$ is the only source of $T_{i}$ for each $i\in\llbracket r\rrbracket$. Let $T$ be the directed tree obtained from $T_{1},\cdots,T_{r}$ by adding a new vertex $z$ and directed edges $(z,z_{i})$ for each $i\in\llbracket r\rrbracket$, and let $\omega$ be the mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ defined as follows: $\omega(z)=0$ and $\omega(v)=\omega_{i}(v)$ whenever $v\in V(T_{i})$, where $i\in\llbracket r\rrbracket$. By the definition of $T$ and $\omega$, it can be verified easily that $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$. Clearly, $r\geq\Delta(T)\geq od_{T}(z)=r$, implying that $\Delta(T)=r$. By Lemma 3.2, $\lambda(T)=\alpha$. $\Box$
@@
Corollary 3.4.
Assume that $0\in{\mathscr{W}}_{r}(\alpha)$. If ${\mathscr{W}}_{r,\alpha}(0)$ contains $\beta$ such that $\beta=\alpha-\sum\limits_{i=1}^{r-1}q_{i}^{-1}>0$ holds for a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r-1\rrbracket\big{\}}$, then there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$.
4 Proof of Theorem 1.5

### v3

# Existence of trees with prescribed maximum degrees and spectral radii

@@ Abstract
It is well known that the spectral radius $\rho(T)$ of a tree $T$ with at least $3$ vertices has the property that $\frac{1}{4}\rho(T)^{2}+1<\Delta(T)\leq\rho(T)^{2}$, where $\Delta(T)$ is the maximum degree of $T$. Let $\mathbb{P}$ denote the set of spectral radii of all non-trivial trees. In this article, we study the inverse problem that for any $\alpha\in\mathbb{P}$ and integer $r$ satisfying the condition $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, is there a tree $T$ such that $\Delta(T)=r$ and $\rho(T)=\alpha$?
For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote a set of non-negative real numbers defined as follows: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,1\leq i\leq s\}$, if $\beta:=\alpha-\sum\limits_{i=1}^{s}q_{i}^{-1}\geq 0$ and $s\leq r-\left\lceil\frac{\beta}{\beta+1}\right\rceil$, then $\beta\in{\mathscr{W}}_{r}(\alpha)$. We first show that $0\in{\mathscr{W}}_{r}(\alpha)$ if and only if there exists a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$. It follows directly that $\mathbb{P}$ is exactly the set of positive numbers $\alpha$ such that $0\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$. Applying this conclusion, we prove that for any two positive integers $r\geq 2$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\rho(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
@@ 1 Introduction
All graphs in this article are finite, simple and undirected. For a graph $G$, let $V(G)$, $E(G)$ $\Delta(G)$ denote its vertex set, edge set and maximum degree, respectively. If $|V(G)|=n$, then $G$ has exactly $n$ eigenvalues, and they are all real numbers (see [6]). The largest eigenvalue of $G$, denoted by $\rho(G)$, is called the spectral radius of $G$.
@@
The spectral radii of graphs, which have been studied extensively (see a survey [33] due to Stevanovíc), plays a crucial role in spectral graph theory, with applications in spanning combinatorics, network theory, and even theoretical physics (see [1, 2, 3, 4, 11, 18, 19, 29, 30, 33]). The distribution of spectral radii among different graphs is a topic of ongoing research with connections to various graph properties and invariants (see [7, 9, 16, 18, 19, 23, 27, 29, 30]).
@@
In this article, we focus on the study of spectral radii of trees. There are various known results on the relation between $\rho(T)$ and other parameters of a tree $T$. Lovász and Pelikán [22] determined the maximal and minimal value of $\rho(T)$ over all trees $T$ with a fixed order, Guo and Tan [13] found an upper bound of $\rho(T)$ in terms of the order and matching number of $T$, and Simić and Tos̆ić [31] confirmed the tree that has the largest spectral radius over all trees with given order and maximum degree. Stevanović [32] provided a relation between $\rho(T)$ and $\Delta(T)$, as stated below.
Theorem 1 ([32]).
For any tree $T$ with $|V(T)|\geq 3$, $\frac{1}{4}\rho(T)^{2}+1<\Delta(T)\leq\rho(T)^{2}.$
A graph $G$ is said to be non-trivial if $|V(G)|\geq 2$. Let $\mathbb{P}$ denote the set of $\rho(T)$ over all non-trivial trees $T$. In this article, we will study the following inverse problem:
Problem 1.
Is it true that for any $\alpha\in\mathbb{P}$ and any integer $r$ satisfying the condition $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$, there exists a tree $T$ with $\rho(T)=\alpha$ and $\Delta(T)=r$?
The above inverse problem connects to characterizing possible spectra of graphs and is similar in spirit to the inverse eigenvalue problem for general matrices. Studying inverse problems in graph theory is important because it offers a reverse perspective on fundamental questions, connects structure to invariants, underpins classical existence theorems, inspires constructive algorithms, etc. In fact, many famous problems in graph theory are inverse in nature. For example, the reconstruction conjecture in graph theory (see [34]) says that any connected graph with at least three vertices is determined uniquely by its subgraphs obtained by removing one vertex, the realization problem of degree sequences studies sequences of positive integers which are degree sequences of graphs (see Erdős and Gallai [10], Havel [15] and Hakimi [14]), the inverse spectral problem studies possible sets of real numbers which are spectra of graphs (see [17]), and the inverse Wiener problem studies whether a given integer can be the Wiener index of some graph (see [20, 35, 36]).
Let $\mathbb{N}$$^{+}$ and $\mathbb{R}$$^{+}$ denote the set of positive integers and the set of positive real numbers, respectively. For any $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of non-negative real numbers defined below:
@@
It can be verified that ${\mathscr{W}}_{1}(1)=\{0,1\}$, ${\mathscr{W}}_{1}(\alpha)=\{\alpha\}$ for all $\alpha\in{\mathbb{R}}^{+}$ with $\alpha\neq 1$, ${\mathscr{W}}_{2}(\sqrt{2})=\left\{0,\sqrt{2},(\sqrt{2})^{-1}\right\}$ and ${\mathscr{W}}_{2}(2)=\big{\{}\frac{i+1}{i}:i\geq 1\big{\}}$. Clearly, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ holds for any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$.
Our first aim in this article is to establish the following result, providing a connection between the existence of a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$ and whether $0\in{\mathscr{W}}_{r}(\alpha)$.
Theorem 2.
For any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$, there exists a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$.
A characterization of the set $\mathbb{P}$ follows directly from Theorems 1 and 2.
Corollary 1.
$\mathbb{P}=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)\big{\}}$.
@@
Corollary 2.
$\mathbb{P}=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}(\alpha)\big{\}}$.
Our second aim in this article is to apply Theorem 2 to prove that Problem 1 has a positive answer for the case $\alpha=\sqrt{k}$ for any $k\in{\mathbb{N}}^{+}\setminus\{1\}$. Certainly, it includes the case $\alpha=k\in{\mathbb{N}}^{+}\setminus\{1\}$.
Theorem 3.
For any $k,r\in{\mathbb{N}}^{+}\setminus\{1\}$, there exists a tree $T$ with $\rho(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
The remainder of this article is structured as follows. In Sections 2 and 3, we prove Theorem 2's necessity and sufficiency, respectively. In Section 4, we complete the proof of Theorem 3. Finally, in Section 5, we provide some open problems arising from our work.
2 The necessity of Theorem 2
The Perron–Frobenius theorem, proved by Oskar Perron [26] in 1907 and Georg Frobenius [12] in 1912, is an important result in the study of eigenvalues and eigenvectors of square matrices. Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectral radius $\rho(G)$ of a connected graph $G$ is obtained. Let $\lambda_{i}(G)$ denote the $i$-th largest eigenvalue of $G$. Then, $\rho(G)=\lambda_{1}(G)$.
Theorem 4 (Perron–Frobenius theorem).
@@ 2 The necessity of Theorem 2
Applying Theorem 4, the next conclusion follows directly.
Proposition 1.
Let $G$ be a connected non-trivial graph and $\alpha$ be a positive number. Then, $\rho(G)=\alpha$ if and only if there is a mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
@@
Note that for any mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$, the condition of (1) holds if and only if $\phi$ is an eigenvector of $A(G)$ with respect to $\rho(G)$. Now we are going to prove the following lemma by which the necessity of Theorem 2 follows directly.
Lemma 1.
Let $T$ be a non-trivial tree and $\alpha=\rho(T)$, and let $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ be an eigenvector of $A(T)$ with respect to $\alpha$. Then both $\frac{\phi(u_{1})}{\phi(u_{2})}$ and $\frac{\phi(u_{2})}{\phi(u_{1})}$ belong to ${\mathscr{W}}_{r}(\alpha)$ for each $u_{1}u_{2}\in E(T)$, where $r=\Delta(T)$.
Proof. Since $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is an eigenvector of $A(T)$ with respect to $\alpha=\rho(T)$, (1) holds for each $u\in V(T)$. Let's first prove the following claim.
@@
Proof of the necessity of Theorem 2. We follow the conditions in Lemma 1. Take any edge $u_{1}u_{2}\in E(T)$, where $u_{1}$ is a leaf of $T$. We have $\alpha\phi(u_{1})=\phi(u_{2})$ by (1), implying that $\alpha^{-1}=\frac{\phi(u_{2})}{\phi(u_{1})}\in{\mathscr{W}}_{r}(\alpha)$ by Lemma 1. Thus, $0=\alpha-\frac{1}{\alpha^{-1}}\in{\mathscr{W}}_{r}(\alpha)$ by definition. $\Box$
3 The sufficiency of Theorem 2
@@
By Lemma 2, if $0\in{\mathscr{W}}_{r}(\alpha)$, then ${\mathscr{T}}_{r,\alpha}(0)\neq\emptyset$, and there exists $(T,\omega)$ in ${\mathscr{T}}_{r,\alpha}(0)$.
Lemma 3.
For any $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$, we have $\Delta(T)\leq r$ and $\rho(T)=\alpha$.
Proof. By the definition of ${\mathscr{T}}_{r,\alpha}(0)$, $\Delta(T)\leq r$. It suffices to show that $\rho(T)=\alpha$.
By definition again, $T$ is a directed tree with a unique source $z$ and $\omega(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)$. Let $T_{1},T_{2},\dots,T_{s}$ be the components of TzT-z, where $s=od_{T}(z)$. Clearly, $1\leq s\leq r$. For each $i\in\llbracket s\rrbracket$, $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(\beta_{i})$, where $\omega_{i}$ is the restriction of $\omega$ to $V(T_{i})$, $\beta_{i}=\omega(z_{i})$ and $z_{i}$ is the vertex in $T_{i}$ with $z_{i}\in N^{+}_{T}(z)$. Then $z_{i}$ the unique source of $T_{i}$ for each $i\in\llbracket s\rrbracket$ and
@@
In order to show that $\rho(T)=\alpha$, by Proposition 1, it suffices to prove the following Claim.
@@
where the last step follows from (4). Thus, $\phi$ satisfies condition (1) for each $u\in V(T)$, implying that $\rho(T)=\alpha$ by Proposition 1. $\Box$
By Lemmas 2 and 3, if $0\in{\mathscr{W}}_{r}(\alpha)$, then there exists a tree $T$ with $\Delta(T)\leq r$ and $\rho(T)=\alpha$, and thus the sufficiency of Theorem 2 holds.
Corollary 3.
If there exist a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r\rrbracket\big{\}}$ such that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, then there exists a tree $T$ with $\Delta(T)=r$ and $\rho(T)=\alpha$.
Proof. Assume that $\sum\limits_{1\leq i\leq r}q_{i}^{-1}=\alpha$, where $q_{i}\in{\mathscr{W}}_{r}(\alpha)$ and $q_{i}>0$ for each $i\in\llbracket r\rrbracket$. By Lemma 2, there exists $(T_{i},\omega_{i})\in{\mathscr{T}}_{r,\alpha}(q_{i})$ for each $i\in\llbracket r\rrbracket$. Assume that $z_{i}$ is the only source of $T_{i}$ for each $i\in\llbracket r\rrbracket$. Let $T$ be the directed tree obtained from $T_{1},\cdots,T_{r}$ by adding a new vertex $z$ and directed edges $(z,z_{i})$ for each $i\in\llbracket r\rrbracket$, and let $\omega$ be the mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ defined as follows: $\omega(z)=0$ and $\omega(v)=\omega_{i}(v)$ whenever $v\in V(T_{i})$, where $i\in\llbracket r\rrbracket$. By the definition of $T$ and $\omega$, it can be verified easily that $(T,\omega)\in{\mathscr{T}}_{r,\alpha}(0)$. Clearly, $r\geq\Delta(T)\geq od_{T}(z)=r$, implying that $\Delta(T)=r$. By Lemma 3, $\rho(T)=\alpha$. $\Box$
@@
Corollary 4.
Assume that $0\in{\mathscr{W}}_{r}(\alpha)$. If ${\mathscr{W}}_{r,\alpha}(0)$ contains $\beta$ such that $\beta=\alpha-\sum\limits_{i=1}^{r-1}q_{i}^{-1}>0$ holds for a multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket r-1\rrbracket\big{\}}$, then there exists a tree $T$ with $\Delta(T)=r$ and $\rho(T)=\alpha$.
4 Proof of Theorem 3